# The strong Fe K line and spin of the black-hole X-ray binary MAXI J1631–479

### v1

@@ 4 Discussion
We have obtained a relatively large Eddington ratio, $L_{\rm disk}/L_{\rm E}\approx 0.3$ (see Table 2), confirming the need to use a model of slim disks. The observed bolometric fluxes are 4.1 and $4.5\times 10^{-8}$ erg cm-2 s-1, respectively. However, the source is strongly absorbed, and the unabsorbed bolometric flux is about four times larger, $F_{\rm bol}=1.7$ and $1.8\times 10^{-7}$ erg cm-2 s-1, respectively. The flux in the disk alone (before scattering) is $\approx\!1.4\times 10^{-7}$ erg cm-2 s-1. At the best-fit of the distance and the inclination, the total unabsorbed fluxes correspond to the luminosity of $\approx\!3\times 10^{38}$ erg s-1, assuming the cosine dependence of the flux ($L=2\pi D^{2}F_{\rm bol}/\cos i$).
The reflection fractions are $\approx$2.4 and $\approx$0.8 for Parts I and II at the best fit, respectively, though the former can be significantly lower, ${\cal R}\approx 1.8{+1.2}_{-0.7}$, if the FMP B data are not used. The relatively high values for Part I can be explained by the effect of anisotropy of a slab corona containing relativistic electrons (Ghisellini et al., 1991). The seed blackbody photons from the disk are emitted upward and are scattered by the electrons in the corona. The highest probability of a collision is when electrons are moving downward. When the electrons are relativistic, they scatter preferentially forward, thus in the direction of the disk. This leads to enhancement of the reflection when measured with respect to the radiation emitted to the observer. The enhancement can be large, even by a factor of several (Ghisellini et al., 1991). The higher value for Part I is likely due to its coronal electrons being more relativistic than those of Part II. This is illustrated in Figure 4, which shows the fitted hybrid (Maxwellian followed by a power law) distributions. The hybrid electrons in our calculations have distinct high-energy tails, and a part of them is relativistic, especially for Part I with a large fraction of electrons with $\beta\gamma\gtrsim 1$, as shown in Figure 4. Thus, the anisotropy effect is strong and can easily account for the large reflection fraction (especially when the uncertainty related to the FMP B accuracy is taken into account). Furthermore, the corona for Part I has a much lower optical depth, with $\tau_{\rm T}\approx 0.19$, than $\tau_{\rm T}\approx 0.73$ for Part II. The scattering in Part I is thus mostly single, for which the relativistic kinematics forces it to be largely towards the disk, while $\tau_{\rm T}\sim 1$ for Part II, when multiple scatterings isotropise the emitted flux. Ideally, this effect could be strictly included in comppsc for the slab geometry, in which we could switch between the down and up emissions. A development of such an extension of comppsc is, however, beyond the scope of this work.
Our obtained value of the spin is relatively high, with the fitted $a_{*}\approx 0.8$–0.9. This is much more than most of the spins measured via gravitational waves emitted during mergers of two BHs in binaries, which are $a_{*}\approx 0.1\pm 0.1$ on average (The LIGO Scientific Collaboration et al., 2025). Also, it is more than the typical natal BH spin of $a_{*}\lesssim 0.1$ (Fuller & Ma, 2019; Belczyński et al., 2020). An increase of the spin from $a_{*}=0.1$ to $a_{*}=0.8$ requires past accretion of about 30% of the current BH mass, i.e., $\approx 3{M}_{\sun}$ (using the formulae of Bardeen 1970). This can be achieved only in binary evolution models with the initial masses of donors in BH LMXBs significantly higher than those presently observed (Podsiadlowski et al., 2003; Fragos & McClintock, 2015).
@@
In this work, we modeled both the Comptonization and reflection self-consistently by using convolution models. We assumed that the high-energy tail is from Comptonization of the disk blackbody photons, which appears very likely given the strength of the disk blackbody component. Consequently, the scattered spectrum is strongly curved, with the shape far from a power law, see the green curve in Figure 3. The flux of this spectrum at the Fe K region is a factor of several times higher than the flux of the power-law fitted with reflection to the high-energy tail. Thus, the irradiating flux is enhanced by at least that factor, which, as we found, can explain the presence of the relatively strong Fe K line in the observed spectrum. This clearly demonstrates that spectral fits assuming the incident spectrum being a power law, such as relxill or reflionx, would give incorrect results for this source, as well as for other sources with soft-state high-energy tails. Instead, it is necessary to use convolution models for both Comptonization and reflection.
@@
We are aware of the limitations of our treatment of reflection. The xilconv model is based on the tables used by xillver for the density of $10^{15}$ cm-3, while the disk surface density is likely to be much larger. This can result in an overestimated Fe abundance (García et al., 2018a), as in our case. The limited accuracy of xilconv was pointed out by Ding et al. (2024). Furthermore, the xilver tables have been done for a cold slab heated only by the irradiating photons, whereas in the present case the intrinsic dissipation dominates over heating by irradiation, and the maximum disk blackbody temperature is $kT_{\rm max}\approx 1$ keV (as we have checked replacing slimbh by diskbb (Mitsuda et al., 1984)). However, a more accurate convolution reflection model is not available at this time. Likely, it would have given different values of the fitted ionization parameter, the relative Fe abundance, and the spin. However, we are convinced that our main result, that the observed Fe K line can be explained by the irradiation by the spectrum from Compton scattering of disk blackbody photons, would stand.
@@
Another caveat for our results is that we use the accretion disk model in the region of the Eddington ratio in which the disk is unstable (Lightman & Eardley, 1974; Shakura & Sunyaev, 1976). Accretion disks can be stabilized by large-scale magnetic fields, e.g., Sądowski (2016), Mishra et al. (2020). However, the structure of such disks can be quite different from the standard disks, e.g., Lančová et al. (2019). In particular the ISCO radius may not form a boundary for the flow, e.g., Rule et al. (2025). Still, the constancy of the inner radius in the soft state of BH XRBs found observationally in many sources in the soft state (Gierliński & Done, 2004) provides an argument for an important role played by the ISCO radius.
Our results also confirm the model dependence of the fitted spins of BH XRBs using different methods, see Zdziarski et al. (2025a, b). Here, we obtained $a_{*}=-0.61_{-0.26}^{+0.23}$ using kerrbb and $a_{*}=0.84\pm 0.03$ using slimbh, while both results are different from the two previous measurements giving $a_{*}>0.94$ using the reflection method alone. While we argue that the result with slimbh provides the most accurate results on the grounds of the accuracy of the disk model and the physical plausibility, we are aware of the caveats for it (as discussed above). An independent determination of the mass and the distance of MAXI J1631–479 is highly desirable for testing our model, though it is very difficult given the very high extinction toward the source (Section 1).
5 Conclusions
@@
We have successfully modeled the strong and broad Fe K line in MAXI J1631–479 by reflection of the irradiation of the accretion disk by the spectrum from Comptonization of the disk blackbody by coronal relativistic electrons. We have taken into account the detailed shape of this spectrum, including its curvature reflecting the shape of the disk blackbody, see Figure 3. That flux is significantly higher than that corresponding to a fit with a power-law irradiation and reflection, e.g., using the reflection models relxillCp or reflionx. We have used convolution models for both Comptonization and reflection, and we stress the necessity to use such models in the soft state of BH XRBs.
@@
The distribution of the coronal electrons is hybrid, with a Maxwellian followed by a power-law tail. In Part I of the observation, a significant part of the coronal electrons was relativistic, which led to an enhancement of the flux directed back to the disk with respect to that directed outward (Ghisellini et al., 1991).
@@
In our modeling, we have used two relativistic disk models to measure the BH spin using the continuum fitting method and coupled that with the reflection method. Using the standard continuum model kerrbb, we have obtained a negative spin, $a_{*}=-0.61^{+0.23}_{-0.26}$. Then, we used slimbh, which accounts for both a finite disk scale height at Eddington ratios comparable to unity and for the vertical radiative transfer of the disk emission. That model gave us $a_{*}=0.84^{+0.03}_{-0.03}$, which we consider more likely (though still a significant systematic error) than the former. However, this underlines the strong model dependence inherent to fitting spins of BH XRBs, and we consider that value as tentative. With the slimbh model, we estimated the BH mass as $9.2_{-1.0}^{+1.3}{M}_{\sun}$, the distance as $4.9^{+0.5}_{-0.5}$ kpc, and the disk inclination of $33^{+2}_{-1}\arcdeg$. The source Eddington ratio is $\sim$0.3.
@@ Acknowledgements
We thank Jorge Casares for a discussion on the donor of MAXI J1631–479. We acknowledge support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138. MS acknowledges support from the National Science Center grant 2023/50/A/ST9/00527. BY is supported by the Natural Science Foundation of China (NSFC) grants 12322307, 12361131579, and 12273026, and by the Xiaomi Foundation/Xiaomi Young Talents Program. MP acknowledges support from the JSPS Postdoctoral Fellowship for Research in Japan, grant number P24712, as well as the JSPS Grants-in-Aid for Scientific Research-KAKENHI, grant number J24KF0244.

### v2

@@ IV Discussion
The main goal of this study is explain the strength of the Fe K line, which was found before to be stronger than that expected from a power-law emission of the corona irradiating the disk. In our approach, we accounted for the incident spectra differing from a power law, assuming that the irradiating spectrum results from Comptonization of the disk emission. The Comptonization of disk blackbody photons (as well as reflection) is modeled self-consistently using convolution models. The scattered spectrum is found to be strongly curved, with the shape far from a power law, see the green curve in Figure 3. To quantify this difference, we have fitted a model with slimbh+relxill to the data. relxill calculates relativistic reflection assuming the incident spectrum is an e-folded power law (setting the e-folding energy to 1000 keV). The fluxes of this spectrum at the Fe K edge at 7 keV were found in our model to be a factor of $\approx$2 times higher than the flux of the power-law fitted with reflection to the high-energy tail. Thus, the irradiating flux is enhanced by that factor, which helps explain the relatively strong Fe K line observed in the spectrum. This also demonstrates that spectral fits assuming the incident spectrum is a power law, such as relxill or reflionx, would yield incorrect results for soft-state sources with high-energy tails. Instead, we advocate the use of convolution models.
The fitted reflection fractions are $2.8^{+1.6}_{-1.0}$ and $1.0_{-0.3}^{+0.3}$ for Parts I and II, respectively. This difference is shown in Figure 3, where the reflection spectrum (red) is compared with the Comptonized irradiating spectrum (green). The relatively high values for Part I can then be explained by the effect of anisotropy of a slab corona containing relativistic electrons (Ghisellini et al., 1991). The seed blackbody photons from the disk are emitted upward and are scattered by the electrons in the corona. The highest probability of collision occurs when electrons move downward. When the electrons are relativistic, they scatter preferentially forward, thus in the direction of the disk. This leads to an enhancement of the reflection when measured relative to the radiation emitted toward the observer. The enhancement can be large, even by a factor of several (Ghisellini et al., 1991). The higher value for Part I is likely due to its coronal electrons being more relativistic than those of Part II. This is illustrated in Figure 4, which shows the fitted hybrid (Maxwellian followed by a power law) distributions. The hybrid electrons in our calculations have distinct high-energy tails, and a part of them is relativistic, especially for Part I with a large fraction of electrons with $\beta\gamma\gtrsim 1$, as shown in Figure 4. Thus, the anisotropy effect is strong and can plausibly account for the large reflection fraction.
Furthermore, the corona for Part I has a lower optical depth (though the error bars marginally overlap), with $\tau_{\rm T}=0.25^{+0.30}_{-0.10}$, than $\tau_{\rm T}=0.68^{+0.29}_{-0.18}$ for Part II. The scattering in Part I can be predominantly single, in which case the relativistic kinematics forces it to be largely towards the disk. In contrast, there will be more multiple scatterings in Part II, which isotropise the emitted flux. Ideally, all these effects could be incorporated into comppsc for the slab geometry, where we would switch between down- and up-emission modes. A development of such an extension of comppsc is, however, beyond the scope of this work, where we still use its currently available version. Therefore, while our results demonstrate that the observed strength of the Fe K line can be explained by the two mentioned effects: the curvature of the incident spectrum and the reflection anisotropy, this is not proven owing to our treatment of the latter being only quantitative.
@@
We are aware of the limitations of our treatment of reflection. The xilconv model is based on the tables used by xillver for the density of $10^{15}$ cm-3, while the disk surface density is likely to be much larger. This usually results in a strongly overestimated Fe abundance (García et al., 2018a), as in our case. The limited accuracy of xilconv was pointed out by Ding et al. (2024). Furthermore, the xilver tables have been done for a cold slab heated only by the irradiating photons. In contrast, in the present case, the intrinsic dissipation dominates over heating by irradiation, and the maximum disk blackbody temperature is high, $kT_{\rm max}\approx 1$ keV (which we checked replacing slimbh by diskbb (Mitsuda et al., 1984)). However, a more accurate convolution reflection model is not currently available. It would likely have yielded different values for the fitted ionization parameter, the Fe abundance, and the spin. However, it appears that our main result, that the observed Fe K line fluxes can be explained by the irradiation by the spectrum from Compton scattering of disk blackbody photons, is independent of those details, and it would stand.
The secondary goal is to determine the spin of MAXI J1631–479. In our approach, we followed the method of Parker et al. (2016), in which the continuum and reflection methods are combined, and the distance and BH mass are fitted. However, there is a major limitation to this method when using the kerrbb model. That model employs the thin-disk model of Novikov & Thorne (1973) regardless of the Eddington ratio. Therefore, it gives identical spectra for different combinations of $M$, $D$, and the disk mass accretion rate $\dot{M}_{\rm disk}$ as long as they follow $M\propto D\propto\dot{M}_{\rm disk}^{1/2}$. However, $L_{\rm disk}/L_{\rm E}\propto D^{2}/M$ changes. The fit results then differ depending on the chosen initial values. This is the reason we did not provide these values when reporting the kerrbb results [^5] Therefore, the values of $M$, $D$, and $\dot{M}_{\rm disk}$ found for GX 339–4 by Parker et al. (2016), who used kerrbb are subject to the above scaling, which was pointed out by Parker et al. (2019). The same scaling applies to the kerrbb fits in Zdziarski et al. (2025), where, however, the parameters fitted to other disk models were found preferable. . Still, the spin and inclination can be obtained with this method.
On the other hand, slimbh spectra depend on the Eddington ratio, based on the slim disk model. Furthermore, they include the radiative transfer calculations of Davis et al. (2005); Davis & Hubeny (2006) as implemented in the bhspec model [^6] https://heasarc.gsfc.nasa.gov/docs/software/xspec/models/bhspec.html. Those explicitly depend on both the mass and the Eddington ratio. This breaks the degeneracy present in kerrbb and allows us to estimate the mass and distance.
Our results confirm the model dependence of the fitted spins of BH XRBs, as shown using different methods (see Zdziarski et al. 2025, 2026). Here, we obtained $a_{*}=-0.40^{+0.15}_{-0.23}$ using kerrbb and reflection and $a_{*}=0.86^{+0.03}_{-0.03}$ using slimbh and reflection. At the same time, both results differ from the three previous measurements, yielding $a_{*}>0.94$ using the reflection method alone. Retrograde accretion in a BH XRB is possible but rare. Outside dense clusters, it requires a BH spin reversal after its formation. Zdziarski et al. (2025) reviewed four published claims of a negative spin in an XRB, and found none of them to be fully convincing.
While the results with slimbh appear to provide the currently most accurate estimates, we are aware of their uncertainties. An independent determination of the mass and the distance of MAXI J1631–479 is highly desirable for testing our model. However, it is tough given the very high extinction toward the source (Section I).
Another caveat regarding the spin value is the possible presence of a warm and optically thick corona. Such coronae appear to be common in active galactic nuclei; see, e.g., Petrucci et al. (2020) and Ballantyne et al. (2024). As we found in Section III.3, the best fit including this model yields $a_{*}=0^{+0.08}$, which is similar to the results for a few other BH XRBs (see Section III.3). Physically, the reduction in the fitted spin is due to the lower maximum temperature of the underlying disk (see figure 6b in Zdziarski et al. 2024b), with the remaining hard emission accounted for by the warm corona. The temperature reduction leads to an increase in the model value of the disk’s inner radius and, consequently, a decrease in the fitted spin. While in MAXI J1631–479 there is no statistical preference for this model, its presence cannot be ruled out either.
Our obtained spin value is relatively high, with the fitted $a_{*}\approx 0.8$–0.9. This is much more than most of the spins measured via gravitational waves emitted during mergers of two BHs in binaries, which are $a_{*}\approx 0.1\pm 0.1$ on average (The LIGO Scientific Collaboration et al., 2025). Also, it is more than the typical natal BH spin of $a_{*}\lesssim 0.1$ (Fuller & Ma, 2019; Belczyński et al., 2020). An increase of the spin from $a_{*}=0.1$ to $a_{*}=0.85$ requires past accretion of about 35% of the current BH mass, i.e., $\gtrsim 3.5{M}_{\sun}$ (using the formulae of Bardeen 1970). This can be achieved only in binary evolution models in which donor initial masses in BH LMXBs are significantly higher than those presently observed (Podsiadlowski et al., 2003; Fragos & McClintock, 2015).
@@
Another caveat for our results is that the standard disk model predicts that the disk is viscously and thermally unstable when dominated by radiation pressure (Lightman & Eardley, 1974; Shakura & Sunyaev, 1976). This is contrary to observations of the X-ray emission in the soft states of BH XRBs showing their disks to be stable (Gierliński & Done, 2004), which is also the case for MAXI J1631–479 (see Table 2). Accretion disks can be stabilized by large-scale magnetic fields, e.g., Begelman & Pringle (2007), Sądowski (2016), Mishra et al. (2020). The structure of such disks can differ substantially from that of standard disks, as noted by, e.g., Lančová et al. (2019). In particular, the ISCO radius may not form a boundary for the flow, e.g., Rule et al. (2025). Still, the constancy of the inner radius in the soft state of BH XRBs, observed in many sources (Gierliński & Done, 2004), supports an important role for the ISCO radius.
While the disk is stable, the corona shows strong variability, see Table 2. This behavior of ‘stable disk/unstable corona’ is typical for the soft state of BH XRBs, as first found by Churazov et al. (2001), and later in several sources, e.g., Swift 1727.8–1613 (Chand et al., 2026). It indicates strongly variable energy deposition in the corona, possibly related to magnetic buoyancy or reconnection. In our models, we have considered the average spectra of both the corona emission and the disk reflection. Since they are expected to be strongly correlated, this approach should not introduce a bias.
We have obtained a relatively large Eddington ratio, $L_{\rm disk}/L_{\rm E}\approx 0.25$ (see Table 3), which shows the need to use a slim-disk model. The observed bolometric fluxes are 4.1 and $4.5\times 10^{-8}$ erg cm-2 s-1, respectively. However, the source is strongly absorbed, and the unabsorbed bolometric flux is about four times larger, $F_{\rm bol}=1.7$ and $1.8\times 10^{-7}$ erg cm-2 s-1, respectively. The flux in the disk alone (before scattering) is $\approx\!1.4\times 10^{-7}$ erg cm-2 s-1. At the best-fit of the distance and the inclination, the total unabsorbed fluxes correspond to the luminosity of $\approx\!3.5\times 10^{38}$ erg s-1, assuming the cosine dependence of the flux ($L=2\pi D^{2}F_{\rm bol}/\cos i$).
V Conclusions
@@
We have modeled the strong and broad Fe K line in MAXI J1631–479 by reflection of the irradiation of the accretion disk by the spectrum from Comptonization of the disk blackbody by coronal relativistic electrons. We have taken into account the detailed shape of this spectrum, including the crucial effect of the spectral curvature, which reflects the shape of the disk blackbody (see Figure 3). Then, that flux is twice higher than that corresponding to a fit with a power-law irradiation and reflection, e.g., using the reflection models relxill or reflionx. We have employed convolution models for both Comptonization and reflection, and we emphasize the necessity of using such models in the soft state of BH XRBs.
@@
The distribution of the coronal electrons is hybrid, with a Maxwellian followed by a power-law tail. In Part I of the observation, a significant part of the coronal electrons was relativistic, which then leads to an enhancement of the flux directed back to the disk with respect to that directed outward (Ghisellini et al., 1991). These two effects can plausibly explain the presence of the strong Fe K line, though the enhancement of the downward flux was estimated only qualitatively.
@@
We used two relativistic disk models to measure the BH spin via the continuum-fitting method coupled to the reflection method. Using the standard continuum model kerrbb, we have obtained a negative spin, $a_{*}=-0.40^{+0.15}_{-0.23}$. Then, we used slimbh, which accounts for both a finite disk scale height at Eddington ratios near unity and the vertical radiative transfer of disk emission. That model gave us $a_{*}=0.86^{+0.03}_{-0.03}$, which we consider more likely (though still bearing a significant systematic error) than the former. However, this underscores the strong model dependence inherent in fitting the spins of BH XRBs, and we consider that value tentative. With the slimbh model, we estimated the BH mass as $10.7_{-0.9}^{+1.5}{M}_{\sun}$, the distance as $5.1^{+1.1}_{-0.4}$ kpc, and the disk inclination of $36^{+2}_{-2}\arcdeg$. The source Eddington ratio is $\sim$0.25.
@@ Acknowledgements
We thank Jorge Casares for a discussion on the donor of MAXI J1631–479, and the two referees for valuable comments. We acknowledge support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138. MS acknowledges support from the National Science Center grant 2023/50/A/ST9/00527. BY is supported by the Natural Science Foundation of China (NSFC) grants 12322307, 12361131579, and 12273026, and by the Xiaomi Foundation/Xiaomi Young Talents Program. MP acknowledges support from the JSPS Postdoctoral Fellowship for Research in Japan, grant number P24712, as well as the JSPS Grants-in-Aid for Scientific Research-KAKENHI, grant number J24KF0244.